### v1

# LLM App Store Analysis: A Vision and Roadmap

Abstract.
The rapid growth and popularity of large language model (LLM) app stores have created new opportunities and challenges for researchers, developers, users, and app store managers. As the LLM app ecosystem continues to evolve, it is crucial to understand the current landscape and identify potential areas for future research and development. This paper presents a forward-looking analysis of LLM app stores, focusing on key aspects such as data mining, security risk identification, development assistance, etc. By examining these aspects, we aim to provide a vision for future research directions and highlight the importance of collaboration among stakeholders to address the challenges and opportunities within the LLM app ecosystem. The insights and recommendations provided in this paper serve as a foundation for driving innovation, ensuring responsible development, and creating a thriving, user-centric LLM app landscape.
conference: International Workshop on Software Engineering in 2030; November 2024; Puerto Galinàs (Brazil)
Figure 1. LLM app ecosystem components and operating mechanisms.
1. Introduction
Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), are trained on vast amounts of text data, allowing them to capture the intricacies of language and perform a wide range of natural language processing tasks. The advent of LLMs has opened up new possibilities for various applications, including chatbots, content generation, language translation, and sentiment analysis. As the capabilities of LLMs continue to expand, there has been a growing interest in making these models accessible to a broader audience. This has led to the emergence of LLM app stores, such as OpenAI’s GPT Store (OpenAI, 2024a), Poe (Quora, 2024), and FlowGPT (Wang and Dang, 2024), which provide a platform for developers to showcase their LLM-powered apps and for users to discover and engage with these apps. LLM app stores offer a centralized marketplace where users can browse, download, and use LLM-based apps across various domains, such as productivity, education, entertainment, and personal assistance.
$1.$ Definitions
To provide the foundation for our analysis and discussions, it is essential to establish clear definitions for the key concepts and terms related to LLM app and LLM app store. The following definitions explain these elements, serving as a reference point for the subsequent sections of this paper.
  • LLM app: A specialized application (app) powered by an LLM, distinct from conventional mobile apps that may incorporate LLM technology. These apps, typically found on platforms like OpenAI’s GPT store, Poe, and FlowGPT, are specifically designed to harness the advanced capabilities of LLMs for a variety of purposes, tasks, or scenarios.
  • LLM app store: A centralized platform that hosts, curates, and distributes LLM apps, enabling users to discover and access tailored intelligent services.
As illustrated in Figure 1, the LLM app ecosystem presents a collaborative environment that harnesses the power of LLMs to create tailored AI apps for a wide range of users. In this ecosystem, LLM app store managers play a key role by boosting the visibility and reach of LLM apps. They support LLM app developers by providing essential resources such as comprehensive documentation, technical support, and marketing assistance, facilitating the creation and launch of cutting-edge LLM apps. Additionally, they ensure a user-friendly experience for end-users with easy-to-use search and navigation features, helping them find the LLM apps that best meet their needs. Moreover, these managers smooth out the transaction process, enabling developers to profit from their innovations.
LLM app developers (creators) are the driving force behind the ecosystem’s innovation. They create and customize LLM apps tailored to unique requirements and use cases. During the development process, developers design instructions and define the desired capabilities for their apps, such as web browsing, image generation, or code interpretation. They can further enrich their apps by uploading external knowledge sources or integrating third-party services through API keys, OAuth protocols (Leiba, 2012), or other authentication mechanisms, enhancing the apps’ functionality and versatility. Once developed, LLM app developers can deploy and submit their apps to the platforms, making them accessible to other users. These distribution channels include the official LLM app stores, third-party LLM app stores, as well as social media like Twitter, and Reddit, or search engines like Google, Bing, Yandex, and Baidu, providing multiple avenues for discoverability. End-users, encompassing individuals, businesses, or organizations, form the consumer base of the LLM app ecosystem. They can browse and discover available LLM apps through the various distribution channels, purchase or acquire them, and provide reviews and feedback to help improve the ecosystem.
1.2. Importance of analyzing LLM app stores
The LLM app store ecosystem is experiencing rapid growth and diversification, as evidenced by several key players and milestones in the field. FlowGPT (Wang and Dang, 2024) stands as a prime example of the vast potential within LLM app stores, boasting over 4 million monthly active users. Moreover, it has recently secured a significant milestone by completing a \$10 million Pre-A funding round (Newswire, 2024), underscoring its growing influence and success in the sector. Additionally, OpenAI’s GPT Store (OpenAI, 2024a) leading this evolution by hosting over 3 million apps (OpenAI, 2024b). In the burgeoning third-party LLM app store arena, as of April 1, 2024, the landscape is diverse and expansive: GPTs App (Prosser, 2024) dominates with 801,185 apps, and GPTs Hunter (Hunter, 2024) is not far behind, offering a substantial repository of 519,000 apps. Meanwhile, GPTStore.AI (GPTStore.AI, 2024) provides a solid selection of 179,895 apps, and GPTs Works (Idoubi, 2024) contributes 103,739 apps, each platform adding unique value and perspective to the LLM app ecosystem.
This rapid expansion parallels the earlier trajectories observed in traditional mobile app stores (Harman et al., 2012), where the proliferation of apps necessitated advanced analytical approaches to ensure quality, security, and relevance. Just as mobile app store analysis has become indispensable in optimizing user experience and app performance (Zhu et al., 2024), a similar emphasis on LLM app store analysis is crucial (Zhang et al., 2024). This new domain presents unique challenges and opportunities, from ensuring the ethical deployment of LLM technologies to navigating the complex dynamics of user engagement and content moderation. Unfortunately, the academic landscape in this area remains starkly underexplored, presenting an extensive frontier teeming with opportunities for inquiry.
The burgeoning LLM app ecosystem offers a fertile ground for in-depth exploration. Investigating LLM app stores is pivotal for gaining insights into the dynamics of LLM apps in real-world scenarios, encompassing user engagement, market dynamics, and technological trends. This examination can highlight best practices, pinpoint prevailing challenges, and spotlight areas ripe for enhancement. Furthermore, delving into LLM app stores can illuminate the broader societal impacts of LLM-driven applications. As they gain ubiquity, it becomes imperative to scrutinize their utility, the nature of the content they deliver, and their influence on user choices and behaviors. Analysis of elements such as user feedback, app narratives, and promotional content within these platforms could reveal underlying biases, potential misinformation issues, or privacy concerns linked to LLM apps. This exploration can, therefore, provide critical guidance for developers and policymakers in crafting more ethically aligned and user-centric LLM apps, ensuring that these tools contribute positively to societal progress.
Figure 2. LLM app store mining and analysis roadmap.
1.3. Overview
This paper aims to provide a forward-looking analysis of LLM app stores, focusing on key aspects that shape the user experience, developer strategies, and the dynamics of the ecosystem. Through an exploration of LLM app data, security, privacy, and market dynamics, we aim to uncover trends, pinpoint challenges, and highlight opportunities that could inform future research directions. Rather than proposing a specific framework or solution, this paper serves as a visionary document, highlighting the importance of collaboration and shared responsibility among stakeholders in addressing the challenges and leveraging the opportunities presented by LLM app stores. We believe that by providing insights and recommendations based on a comprehensive analysis of the current landscape, this paper can contribute to the development of a thriving, user-centric, and responsible LLM app ecosystem.
In the following sections, we present the roadmap for mining and analyzing LLM app stores, comprising three key stages as shown in Figure 2. The data collection stage (§2) involves gathering and preprocessing LLM app raw data, metadata, and user feedback from LLM app stores. The security and privacy analysis stage (§3) focuses on identifying potential risks and regulatory compliance issues. The ecosystem and market analysis stage (§4) leverages the collected data to gain insights into developer engagement, market trends, and strategic decision-making within the LLM app ecosystem. §5 discusses the implications of our analysis, challenges faced by the ecosystem, and recommendations for stakeholders. Finally, §6 concludes the whole paper.
2. Data collection and preprocessing
To conduct a comprehensive analysis of LLM app stores, researchers must identify the key data types to collect and preprocess. This section outlines the essential data categories for understanding the LLM app ecosystem, including LLM app raw data, LLM app metadata, and user feedback, as illustrated in Figure 2. The following subsections will delve into each data category, followed by a discussion on the importance of data preprocessing in ensuring data quality and preparing the dataset for analysis.
2.1. LLM app raw data
LLM app raw data encompasses various components that define the behavior and capabilities of the LLM apps. Instructions play a vital role in specifying the desired functionality and behavior of the app, outlining actions to perform and those to avoid. Knowledge files provide custom information that the LLM app can access to inform its responses, retrieving relevant sections based on user input. These files may be viewable by other users through LLM app responses or citations, enhancing transparency and trust.
Authentication mechanisms, such as API keys or OAuth protocols (Leiba, 2012), are necessary when LLM apps require integration with third-party services, ensuring secure access. Additionally, LLM apps must adhere to the privacy policies of any integrated third-party platforms to maintain user confidentiality.
Conversation starters are designed to guide new users in asking better questions, providing a smooth onboarding experience. Lastly, custom temperature settings allow for controlling the creativity of the LLM app’s responses, balancing variation and predictability to suit different use cases.
$2.$ LLM app metadata
LLM app metadata plays a crucial role in helping users navigate the LLM app store, providing essential information about each app to facilitate discovery, understanding, and comparison. The app name and creator are fundamental pieces of metadata, allowing users to identify and attribute each app to its respective developer. A detailed description of the app’s purpose and features is essential for users to grasp the app’s intended use case and capabilities quickly.
Capabilities provide users with a clear understanding of the app’s functionalities. They can include a wide range of features, such as web browsing, which enables the app to access and retrieve information from the internet; image generation, allowing users to create visual content through the app; and code interpretation, enabling the app to understand and execute programming languages. Other potential capabilities include speech generation, video generation, etc.
Categories group LLM apps by their primary function or domain, such as productivity, entertainment, or education. Tags provide more granular information about the app’s features, use cases, or compatibility. For example, tags may indicate whether an app is suitable for beginners, works offline, or integrates with specific platforms. The updated time informs users about the app’s currency, ensuring they can access the latest features and content.
Sample chats showcase the app’s conversational abilities, response quality, and potential use cases, giving users a realistic preview of what to expect. Frequently Asked Questions (FAQs) constitute a critical component that systematically addresses prevalent user inquiries. It provides quick answers to common queries about the app’s functionality, limitations, and best practices.
2.3. User feedback
User feedback is a valuable source of data for assessing the performance and popularity of LLM apps. One of the key metrics is the number of conversations, which indicates the level of user engagement with the app. A high number of conversations suggests that users find the app valuable and engaging, regularly interacting with it to fulfill their needs. The retention rate measures the percentage of users who continue to use the app over a specific period. Daily active users (DAU) provide a snapshot of the app’s active user base, representing the number of unique users who engage with the app daily. Tracking DAU over time offers insights into the app’s ongoing appeal and growth trajectory.
Ratings and the number of ratings offer a quantitative measure of user satisfaction, allowing users to express their opinions on a standardized scale. A high average rating and a large number of ratings signify that users generally have a positive experience with the app and are willing to share their feedback. Rankings provide a comparative measure of an app’s performance against other similar apps within the store.
User reviews offer qualitative feedback, allowing users to share detailed opinions, experiences, and suggestions. Positive reviews highlight an app’s strengths and the value it provides to users, while negative reviews can reveal weaknesses, bugs, or areas for improvement. Analyzing user reviews can help developers prioritize updates, fix issues, and enhance features based on user preferences. Information about reviewers, such as their user profile or history with the app, can provide additional context and credibility to their feedback. Social media mentions capture an LLM app’s broader impact and popularity beyond the confines of the LLM app store. Users may share their experiences, recommend the app to others, or engage in discussions related to the app on various social media platforms.
2.4. Data preprocessing
Once the data is collected, it must undergo a rigorous preprocessing phase to ensure its quality, security, and compliance with privacy regulations. Preprocessing steps should be applied to ensure data quality and consistency (Alasadi and Bhaya, 2017; García et al., 2015; Mishra et al., 2020). This involves removing duplicate entries, handling missing values, and normalizing text data. Text preprocessing techniques such as tokenization, lowercasing, and removing stop words and punctuation should be employed. Data cleaning steps, such as removing irrelevant or spam reviews and filtering out apps with insufficient information or user engagement, are also necessary. This preprocessing phase is crucial for obtaining high-quality, reliable data for analysis. Furthermore, this phase may involve filtering out sensitive or personal information, removing malicious content, and ensuring adherence to established policies and guidelines. Data normalization and formatting procedures are also applied to facilitate efficient storage, retrieval, and analysis of the collected information. By following this comprehensive data collection and preprocessing approach, researchers can gain a holistic understanding of the LLM app ecosystem, enabling them to conduct in-depth analyses, identify trends and patterns, and ultimately contribute to the advancement and growth of this rapidly evolving field.
3. Security and privacy analysis
As shown in Figure 2, in the evolving landscape of LLM app stores, security and privacy emerge as paramount concerns, necessitating a comprehensive and multifaceted analysis to ensure the integrity and trustworthiness of the ecosystem.
3.1. Security risks
LLM app raw data-related risks. App cloning, where someone unauthorized copies a legitimate app, infringes on intellectual property rights, and potentially introduces security threats or subpar user experiences. In the mobile app ecosystem, app cloning has been a persistent issue, and app stores have employed techniques like code signing (Cooper et al., 2018; Cho et al., 2012) and similarity analysis (Al-Subaihin et al., 2019; Gonzalez et al., 2015) to detect and prevent cloned apps. For LLM app stores, researchers should explore adapting these existing methods, such as code fingerprinting and advanced similarity analysis techniques tailored to LLM apps, to combat app cloning effectively.
App vulnerabilities refer to security weaknesses within LLM apps that attackers can exploit, potentially leading to data breaches or unauthorized activities. For example, these vulnerabilities could arise from inadequate input validation, allowing attackers to perform injection attacks with crafted inputs to elicit unintended responses. This can manipulate LLMs into generating sensitive information, violating content policies, or executing unauthorized actions, compromising app integrity and user data. Additionally, insufficient input checks may render apps vulnerable to jailbreaking, enabling LLMs to output content or perform tasks against terms of service or regulations, raising legal and ethical concerns. These security flaws are also frequently the result of substandard development practices, such as insecure data storage, where sensitive information is poorly protected, making it accessible to unauthorized parties. Weak encryption methods or a lack of robust database security can further exacerbate these issues. Moreover, inadequate authentication mechanisms, including predictable passwords or the absence of multi-factor authentication, can simplify unauthorized access to app functionalities and data. App vulnerabilities are not uncommon in the mobile app ecosystem (Mutchler et al., 2015; Watanabe et al., 2017), with a wealth of established detection techniques available (Abdullah and Zeebaree, 2021; Chen et al., 2022; Shezan et al., 2017). Accordingly, one of the future research directions should be developing tailored solutions to identify and mitigate vulnerabilities specific to LLM apps. This may involve techniques for securing input validation, preventing jailbreaking, enforcing robust authentication, and ensuring secure data storage and transmission within the LLM app ecosystem.
Malicious apps pose a substantial risk in LLM app stores. The malice can manifest in several ways. For example, developers may create LLM apps using instructions or knowledge files containing malicious content, resulting in the app’s knowledge base being tainted with harmful information. Moreover, LLM apps themselves may output content lacking proper constraints, including pornographic or gambling-related information, or even links directing users to malicious websites. Another example is low description-to-behavior fidelity, that is when the actual performance or actions of an app diverge significantly from its documented descriptions or expected behaviors. These phenomena are also prevalent in the mobile app ecosystem. Various tools and techniques have been developed to detect and mitigate malicious mobile apps, such as static and dynamic analysis (Martinelli et al., 2016; Quan et al., 2014; Seo et al., 2014), machine learning-based malware detection (Li et al., 2018; Soni et al., 2020; Wei et al., 2017), and app vetting processes (Backes et al., 2014; Zhang et al., 2013). The unique challenges posed by malicious LLM apps necessitate the development of tailored detection and mitigation strategies. Researchers should focus on developing novel techniques specifically designed to identify and address the distinct threats posed by malicious LLM apps, ensuring a safe and trustworthy ecosystem.
Third-party service integration is another area of concern, as integrating external services or APIs into an app can introduce vulnerabilities or data privacy issues. For example, if the third-party service provider experiences a data breach or has weak security measures, it could compromise the security and privacy of the LLM app and its users. In the mobile app domain, various mature methods have been proposed to address similar issues, including extensive research on third-party library analysis (Feal et al., 2021; Utz et al., 2022; He et al., 2020; Kawabata et al., 2013). To effectively mitigate the risks associated with third-party service integration in LLM apps, developers should adhere to the principle of least privilege, granting only the minimum necessary permissions and access required for the service to function within the app. Robust authentication and authorization mechanisms should be implemented to ensure that only authorized users and processes can interact with the integrated services. Furthermore, encrypting sensitive data both in transit and at rest is crucial when exchanging information with third-party services to protect the confidentiality and integrity of the data. Regular monitoring and auditing of third-party services should also be conducted to detect any suspicious activities or changes in their security posture.
User tracking and profiling without proper consent is another risk, where excessive tracking of user data, behavior, or activities occurs, often for targeted advertising or analyzing user preferences. This can manifest in various harmful ways, such as identity theft, personalized phishing attacks, or unwanted exposure to tailored yet intrusive advertising (Li et al., 2023; Wei et al., 2012; Yalcin, 2017). Moreover, the accumulation and analysis of such data could result in biased or discriminatory outcomes, where decisions made by these LLM apps might favor or disfavor individuals based on their profiled characteristics. This not only undermines user trust but also raises ethical concerns about the fairness and transparency of LLM-powered apps. To mitigate risks associated with user tracking and profiling, LLM app stores should enforce strict privacy policies, obtain explicit user consent, and employ privacy-preserving techniques like differential privacy and data anonymization. Strategies such as regular audits should be adopted to ensure fairness, accountability, and transparency in LLM app decision-making.
Similar to mobile app protection techniques that often involve obfuscation (Continella et al., 2017; Wang et al., 2018), encryption (Silva et al., 2013; Tysowski and Hasan, 2013), and packing (Chen et al., 2018), LLM apps may employ comparable app protection techniques to safeguard their models and data. For example, a potential risk arises from the reliance on third-party frameworks for app protection. To safeguard against model stealing (Tramèr et al., 2016; Orekondy et al., 2019) and unauthorized model reuse (Ji et al., 2018), developers might obfuscate their LLM apps to protect the model itself. However, this obfuscation process could introduce new security risks. It might inadvertently obscure crucial monitoring and debugging features, making it harder to identify and respond to genuine security threats. Additionally, the complexity added by obfuscation could lead to performance degradation, not only affecting the user experience but also potentially introducing vulnerabilities that attackers could exploit.
Advertisement fraud can occur during the user’s interaction with the LLM app, involving deceptive or misleading ad practices, such as hidden payments, unauthorized data collection, or intrusive ad experiences. Mobile app stores have employed ad network monitoring (Haddadi et al., 2010), real-time ad analysis (Shao et al., 2018), and user feedback analysis (Gui et al., 2015) to combat advertisement fraud. For LLM app stores, researchers should explore adapting these techniques and developing new methods tailored to the unique challenges of LLM apps, ensuring a transparent and trustworthy advertising ecosystem.
Market policy violations, where LLM apps breach the LLM app store’s terms of service, content policies, or other regulations governing app publication and monetization, can undermine the LLM app store’s integrity and user trust. Mobile app stores have implemented automated policy compliance checks (Liu et al., 2021; Zimmeck et al., 2016) and app vetting processes (Zhang et al., 2013) to enforce market regulations. In the context of LLM app stores, researchers should focus on developing automated policy compliance checks tailored to the unique characteristics and challenges of LLM apps, ensuring a secure and trustworthy LLM app ecosystem.
LLM app metadata-related risks. Fake apps, designed to impersonate legitimate LLM apps and deceive users or steal sensitive information, pose a significant risk to users. Mobile app stores have implemented app vetting processes (Zhang et al., 2013) and leveraged techniques like app analysis (Rani et al., 2023) and user feedback (Martens and Maalej, 2019) monitoring to identify fake apps. In the context of LLM app stores, researchers should investigate developing advanced natural language processing and multimedia analysis methods to aid in the detection of fake LLM apps, ensuring user safety and trust.
User feedback-related risks. In the context of LLM app stores, security risks should be carefully considered and addressed. One significant risk is ranking fraud, where attackers attempt to manipulate the LLM app store rankings through illegal methods, such as using bot programs to generate fake ratings, downloads, or reviews, or engaging in keyword stuffing. Similar to the mobile app market, where researchers have proposed systems to detect ranking fraud by analyzing leading sessions, rating patterns, and review behaviors (Zhu et al., 2013, 2014), app stores may need to employ advanced techniques to identify and mitigate fraudulent activities aimed at artificially inflating app rankings and popularity.
Another concern is malicious ASO (i.e., App Store Optimization), where attackers exploit irregular methods to falsify user feedback, such as user engagement metrics or app ratings, to artificially boost an app’s search result rankings and discoverability, ultimately gaining higher exposure and usage when users search for related keywords. This issue is analogous to the collusive promotion groups in the mobile app ecosystem, where developers pay service providers to organize groups of attackers to post fraudulent reviews, inflate download numbers, or manipulate app ratings in an attempt to boost their app’s ranking and visibility (Chen et al., 2017; Padilla-Piernas et al., 2019; Wang et al., 2019), which can ultimately undermine the integrity of the app store’s ecosystem if not addressed.
Spam reviews in LLM app stores can also contain malicious content or involve large-scale fake reviews manipulated by bots or manual efforts, intending to inflate the app’s reputation artificially. This issue is well-documented in the mobile app industry, where spam reviews and review fraud have been a persistent challenge. Various detection methods have been established in the mobile app domain (Genc, 2019; Seneviratne et al., 2015). Similarly, LLM app stores must adopt and refine such techniques to preserve the authenticity and reliability of their review mechanisms.
3.2. Privacy protection
Protecting privacy is a critical aspect of LLM app stores. For developers, it is essential to filter out any privacy data that may be included in the instructions or knowledge base files provided to the LLM app. This includes not only personal identifiable information (PII) (Schwartz and Solove, 2011) such as addresses, contact details, and other sensitive data that could compromise individual user privacy, but also extends to sensitive information related to businesses, governmental bodies, and other entities. Protecting this wider range of data ensures the privacy and security of related stakeholders, safeguarding against potential misuse, data breaches, or other forms of exploitation that could have far-reaching consequences. This is similar to the principles and practices adopted in the mobile app industry, where developers are required to implement appropriate data protection measures to safeguard user privacy and comply with relevant regulations, such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017).
Furthermore, developers must also comply with the LLM app store’s privacy policies and relevant legal regulations when collecting user information for personalized fine-tuning or optimization of the app. This involves clearly informing users about the purpose, scope, and manner of data collection and usage, and obtaining user consent. The LLM app store should review the app to ensure compliance with these requirements. Again, this aligns with the standard practices in mobile app stores, where apps are vetted for their data collection and privacy practices, and users are provided with clear information about how their data is being used (Almuhimedi et al., 2015; Shklovski et al., 2014).
From the user’s perspective, privacy data filtering is crucial when interacting with LLM apps. Users’ input may contain private information, and the app should have filtering mechanisms in place to identify and remove this sensitive data, preventing it from being leaked to developers or stored in the knowledge base. This is analogous to the privacy protection measures implemented in mobile apps, where user inputs and data are often processed locally on the device or through secure channels to protect user privacy (Liu et al., 2014; Nema et al., 2022). Additionally, users can expect LLM app stores to provide transparent information about the privacy practices of listed apps, similar to how mobile app stores provide privacy labels and summaries to help users make informed decisions (Apple, 2024b).
4. Ecosystem and market analysis
In the dynamic ecosystem of LLM app stores, the interplay between developer engagement, competitive landscape, and market trends drives innovation and growth. As displayed in Figure 2, developer support mechanisms, strategies for navigating competitive pressures, and responsiveness to evolving market dynamics are crucial for cultivating a vibrant, sustainable marketplace that caters to diverse user needs and preferences while fostering technological advancements.
4.1. Developer engagement
Enhancing support for LLM app developers is essential to fostering a thriving ecosystem. Implementing effective requirements engineering processes and tools can help developers gain clarity on app specifications and functionalities. Although mobile app development benefits from established practices like user story mapping (Cohn, 2004) and wireframing (Brown, 2010), the LLM app ecosystem should develop specialized tools that cater to the unique needs of conversational AI, such as dialogue flow designers, intent mappers, and entity recognizers.
Providing comprehensive development assistance, including documentation, examples, and best practices, can lower entry barriers and guide developers in creating high-quality LLM apps. Drawing inspiration from the mobile app domain’s extensive resources, such as Apple’s Human Interface Guidelines (Apple, 2021b) and Google’s Material Design (Google, 2021b), the LLM app ecosystem should create similar guides tailored to conversational AI, covering topics like prompt engineering (Arvidsson and Axell, 2023), context management (Ge et al., 2023), and multi-turn dialogue handling (Wu et al., 2023).
Offering robust analysis and testing tools or frameworks can assist developers in evaluating app performance, identifying vulnerabilities, and optimizing the user experience, ensuring high-quality output. Mobile app development has benefited from tools like Appium (Verma, 2017) and Espresso (Google, 2021a), which have revolutionized automated testing, enabling developers to catch bugs early and ensure app stability. Similarly, the LLM app ecosystem needs to invest in developing testing frameworks that can simulate user interactions and detect potential biases or inconsistencies in the generated output.
Standardized third-party service interfaces can simplify the integration process for developers. The LLM app store can provide a list of certified service providers or establish partnerships with leading companies in LLMs and knowledge bases, similar to how mobile app stores have streamlined integration with payment gateways (Isaac and Zeadally, 2012; Yang et al., 2019) and analytics (Han et al., 2016; Minelli and Lanza, 2013) providers.
Cross-platform migration tools and support can help developers deploy LLM apps across multiple platforms. In the mobile app development domain, frameworks like React Native (Eisenman, 2015) and Flutter (Kuzmin et al., 2020; Wu, 2018) have greatly simplified the process of building cross-platform apps. The LLM app ecosystem could explore similar solutions that allow developers to write once and deploy across various conversational AI platforms.
Implementing a comprehensive bonus system that rewards developers based on app quality and user feedback can incentivize continuous optimization. Similar to Apple’s app store small business program (Apple, 2021a) offering financial incentives and recognition for high-performing developers, the LLM app ecosystem should consider initiatives that encourage innovation and user satisfaction.
4.2. Competitive landscape
Leveraging user preferences, history, and ratings, LLM app stores can develop sophisticated recommendation algorithms to suggest potentially interesting LLM apps to users. This not only improves app discoverability but also increases user satisfaction and engagement. Mobile app stores have successfully implemented such recommendation systems, with examples like Apple’s app store featuring “Apps You Might Like” and Google Play’s “Recommended for You” sections (Tu et al., 2019; Hwang et al., 2019). However, in the context of LLM app stores, academic research on tailoring recommendation algorithms to this novel domain remains largely unexplored.
To help developers effectively promote their LLM apps, LLM app stores can offer a range of promotion tools and channels. Mobile app stores usually use ads, featured spots, and events to promote LLM apps, leveraging influencer partnerships for added trust and visibility (Lee et al., 2018; Rahman et al., 2019). Similarly, LLM app stores could also include advertising placement options, such as sponsored search results or featured app listings, allowing developers to increase their app’s visibility to potential users. Additionally, LLM app stores can provide promotional opportunities through curated collections, themed showcases, or developer spotlights, highlighting noteworthy LLM apps and their creators.
Curating and featuring top LLM apps is a strategic move by LLM app stores to influence the app market landscape (Hunter, 2024; OpenAI, 2024a; GPTStore.AI, 2024). Through spotlighting apps that excel in innovation and quality, LLM app stores could establish benchmarks and inspire developers to aim high. This not only guides users to superior apps but also rewards developers for outstanding user experiences.
To improve LLM app discoverability, LLM app stores are encouraged to deploy a coherent app classification system. Sorting LLM apps by their functionality, usage scenarios, industries, or target audiences simplifies the search process for users. This not only elevates the user experience but also supports developers in strategically showcasing their apps. For instance, renowned stores like Google Play (Fu et al., 2013; Google, 2024), Apple App Store (Apple, 2024a; Jia et al., 2020), and Blackberry World App Store (Harman et al., 2012) typically employ a consolidated category structure.
4.3. Market trends and innovations
Establishing a trend analysis mechanism is crucial for LLM app stores to uncover and predict LLM app market trends by mining user behavior data, download volumes, and reviews. This helps stores and developers formulate future strategies, such as identifying increasingly popular features or scenarios. LLM app stores can draw inspiration from the well-established practices in traditional mobile app stores, which have successfully employed trend analysis techniques to identify emerging app trends and user preferences (Fu et al., 2013; Liu et al., 2015; Milward et al., 2016). For instance, analyzing in-app user behavior patterns, feature usage, and navigation paths can provide valuable insights into user preferences and emerging trends for LLM apps.
Analyzing and comparing consumer preferences for LLM apps across different regional markets and cultural backgrounds can reveal market differences, enabling LLM app stores to adjust product strategies and operational approaches accordingly. Cross-cultural research on mobile app adoption has highlighted the importance of tailoring app interfaces, content, and functionality to cater to diverse cultural norms and expectations, which can significantly impact user engagement and retention (Light et al., 2018; Sultan et al., 2009). LLM app stores can leverage these learnings and adapt their offerings, marketing strategies, and localization efforts to better resonate with users from various cultural backgrounds.
User review analysis is a vital channel for LLM app stores to understand genuine user feedback and identify areas for improvement. By applying natural language processing and sentiment analysis to a vast number of user reviews, stores can gain insights into user pain points, app deficiencies, bugs, user expectations, suggestions, and overall acceptance and trust levels for LLM apps. Just as in the mobile app domain, where user reviews have been extensively leveraged to improve app quality and user experience (Guzman and Maalej, 2014), LLM app stores can benefit from similar techniques to mine valuable feedback from user reviews.
As artificial intelligence apps, the development of LLM apps must adhere to human ethical values and maintain a high degree of alignment with humanistic ideals, which is essential for gaining public trust and recognition. App stores should establish review standards that prohibit the listing of LLM apps containing content that violates social morality or harms public interests. The design of LLM apps should embody human-centric values, such as respect for privacy, explainability, and controllability. The functions, algorithms, knowledge bases, and other aspects of LLM apps must align with human interests and avoid producing harmful effects. Similar to the guidelines and best practices established in the mobile app industry for protecting user privacy, ensuring data security, and promoting ethical app development (Balebako et al., 2014; Martínez-Pérez et al., 2015), LLM app stores can adopt and adapt these principles to address the unique challenges and risks associated with AI-powered apps.
5. Discussion
The analysis of LLM app stores following the proposed roadmap has several implications for the development and regulation of these platforms. In addition, this section also discusses the key challenges and provides recommendations for LLM app store stakeholders.
5.1. Implications
The burgeoning LLM app store ecosystem presents a unique blend of opportunities and challenges. For developers, the potential to create innovative, AI-driven apps is immense. Yet, this potential comes with the responsibility to ensure that apps are secure, privacy-compliant, and ethically aligned. The detailed analysis of LLM app raw data, metadata, and user feedback is crucial for developers to understand user needs and preferences, enabling them to design more engaging and useful apps.
For regulators and LLM app store managers, the rapid evolution of LLM app stores necessitates a proactive approach to governance. Ensuring a safe, trustworthy, and inclusive platform requires continuous monitoring for security threats such as malicious apps, spam reviews, and ranking fraud. Furthermore, privacy protection remains paramount, demanding stringent measures to safeguard user data from unauthorized tracking, profiling, and third-party service vulnerabilities.
5.2. Challenges
The analysis of LLM app stores following the proposed roadmap reveals several challenges that need to be addressed to ensure the sustainable growth and responsible development of this ecosystem.
Data privacy and security. The integration of third-party services and the collection of user data by LLM apps raise significant privacy and security concerns. Ensuring compliance with data protection regulations, such as GDPR and CCPA, and implementing robust security measures to prevent data breaches and unauthorized access to user information are critical challenges that require attention from both developers and platform providers.
Intellectual property protection. The prevalence of app cloning and the potential for intellectual property infringement within LLM app stores pose a significant challenge to developers and platform owners. Detecting and preventing the unauthorized copying or reuse of app code, designs, and features is crucial to maintaining a fair and competitive environment that rewards innovation and original work.
Ensuring app quality and reliability. With the rapid growth of LLM app stores, maintaining high standards of app quality and reliability becomes increasingly challenging. Implementing effective app review processes, establishing clear guidelines for developers, and continuously monitoring app performance and user feedback are essential to provide users with a consistent and trustworthy experience.
Addressing algorithmic biases and fairness. LLM apps rely on complex algorithms and models that may inadvertently perpetuate biases or discriminate against certain user groups. Identifying and mitigating these biases, ensuring fairness in app recommendations and search results, and promoting diversity and inclusivity within the app ecosystem are significant challenges that require ongoing research and collaboration between developers, researchers, and LLM app store managers.
Balancing innovation and responsibility. The rapid advancements in LLM technologies and the increasing capabilities of LLM apps present both significant opportunities for innovation and formidable challenges in terms of responsible development and deployment. Striking the right balance between pushing the boundaries of what is possible and considering the ethical, social, and long-term implications of LLM apps is a critical challenge that requires input from multiple stakeholders, including developers, researchers, policymakers, and users.
User education and awareness. As LLM apps become more prevalent and influential in various domains, educating users about their capabilities, limitations, and potential risks becomes increasingly important. Providing clear and accessible information about how LLM apps work, what data they collect, and how users can control their interactions with these apps is a significant challenge that requires collaboration between developers, platform providers, and educational institutions.
Regulatory and policy challenges. The rapid growth and evolving nature of the LLM app ecosystem present challenges for regulatory bodies and policymakers. Developing appropriate legal frameworks, guidelines, and standards that promote innovation while protecting user rights and ensuring accountability is a complex task that requires ongoing dialogue and collaboration between industry stakeholders and policymakers.
5.3. Recommendations for LLM app store stakeholders
For LLM app store managers, it is crucial to implement robust vetting processes for app submissions, incorporating automated and manual review mechanisms to ensure compliance with store policies and ethical standards. Establishing transparent guidelines and providing resources for developers on best practices, security, and privacy can enhance the overall quality of the ecosystem. Furthermore, operators should invest in advanced algorithms for fraud detection and user review analysis to proactively address security risks and improve user trust.
Developers are encouraged to prioritize security and privacy in the design and development of their LLM apps. This includes adhering to best practices for data handling, implementing robust authentication mechanisms, and ensuring transparency in app functionalities and data usage policies. Engaging with the user community to gather feedback and continuously iterating on app features based on user insights can also drive improvement and innovation.
Researchers and policymakers play a pivotal role in shaping the future of LLM app stores. Conducting in-depth studies on user behavior, market trends, and security challenges can provide valuable insights for all stakeholders. Moreover, developing frameworks and guidelines for ethical AI use, privacy protection, and security in LLM apps can guide developers and store operators in creating a responsible and user-centric ecosystem.
For users, staying informed about the apps they use is key. This includes reviewing app permissions, understanding data usage policies, and providing constructive feedback to developers. By actively participating in the ecosystem, users can contribute to the improvement of LLM apps and help foster a culture of transparency and accountability.
The rapidly evolving nature of LLM app stores offers vast potential for innovation and growth. However, realizing this potential requires concerted efforts from all stakeholders to address the challenges of security, privacy, and ethical considerations, ensuring a thriving and sustainable ecosystem for the future.
6. Conclusion
This paper provides a forward-looking analysis of LLM app stores, focusing on key aspects such as app data collection, security and privacy analysis, and ecosystem and market analysis. Through this exploration, we underscore the importance of user-centric design, data privacy, intellectual property protection, and collaboration among stakeholders in shaping the future of the LLM app ecosystem. As the LLM app landscape continues to evolve, ongoing research and collaboration among researchers, developers, LLM app store managers, and policymakers are crucial to address challenges, leverage opportunities, and drive responsible innovation.